Week of 200420

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • remote: Kate (WLCG, chair), Julia (WLCG), Maarten (WLCG, ALICE), Borja (monitoring), Ivan (ATLAS), Jens (NDGF), Roberto (storage), Christoph (CMS), Xin (BNL), Andrew (TRIUMF), Sang Un (KISTI), Mark (LHCb), Dave (FNAL), David (IN2P3), Darren (RAL), Pepe (PIC)

Experiments round table:

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • Preparing for the next Stripping round
      • Ongoing WG MC productions

Sites / Services round table:

  • ASGC: nc
  • BNL: NTR
  • CNAF: NTR
  • EGI: nc
  • FNAL: NTR
  • IN2P3: NTR
  • JINR: All green. The DDM quota for the site may be increased. At the moment we have (at https://test-cmssst.web.cern.ch/cgi-bin/set/SiteCapacity) 9300 TBytes of Disk space usable by CMS.
  • KISTI:
    • Migration of CE from CREAM to HTCondor was completed. Thanks to Maarten for help.
    • A network incident is ongoing since last Friday - while the primary link is under maintenance, the secondary is experiencing asymmetric routing issues
  • KIT: NTR
  • NDGF:
    • During the past weekend and early this week some Atlas and Alice storage is offline and two CE:s MIA due to computer room electricity maintenance in two separate computer rooms.
    • dCache 6.1 upgrade on upcoming Wednesday.
  • NL-T1:
  • NRC-KI: nc
  • OSG: nc
  • PIC:
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services:
    • AIADM now using 2FA on ssh as per OTG:0055962
    • Centos 8.1 cloud images now available
  • CERN storage services: NTR
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • GGUS:146482 Missing answer in a recomputation request for IN2P3-CC from ATLAS and LHCb, which is blocking the final report generation.
  • MW Officer:
  • Networks: NTR
  • Security: NTR

AOB: Julia reminded that validation of both storage and CPU accounting data is pending. Data have to be validated today, otherwise monthly reports (to be sent out tomorrow) will not be correct.

Edit | Attach | Watch | Print version | History: r20 < r19 < r18 < r17 < r16 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r20 - 2020-04-20 - JensHenrikLarsson
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback